OpenStudy (gracygirl):

Write the point-slope form of the equation of the line that passes through the points (6, -9) and (7, 1). Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution. y - y1 = m(x - x1) y - 1 = 1/10(x - 6) y - 1 = 1/10x - 6 y - 1 + 1 = 1/10x - 6 + 1 y = 1/10x - 5 What am I doing wrong?

OpenStudy (gracygirl):

@Will.H @AloneS @Awolflover1 @Frozen15

OpenStudy (studygurl14):

First off, your slope is wrong. Second, why did you change it to slope-intercept form?

OpenStudy (studygurl14):

Slope formula with two points \((x_1,y_1)\) and \((x_2,y_2)\) m=\(\Large\frac{y_2-y_1}{x_2-x_1}\) Point-slope formula: \(\large y-y_1=m(x-x_1)\)

OpenStudy (gracygirl):

I don't know...my teacher didn't even tell me I did that...that's why I'm finally coming here because I'm desperate to get this right. @StudyGurl14

OpenStudy (studygurl14):

Use the two given points and plug into the formula I gave

OpenStudy (gracygirl):

\[m = \frac{ 1 - (-9) }{ 7 - 6 }\]

OpenStudy (gracygirl):

?

OpenStudy (studygurl14):

Yes

OpenStudy (gracygirl):

So that's the answer? Or simplify it?

OpenStudy (gracygirl):

Simplified would be... \[m = \frac{ 1 + 9 }{ 7 - 6 } = \frac{ 10 }{ 1 } = 10?\]

OpenStudy (gracygirl):

@StudyGurl14

OpenStudy (studygurl14):

yes

OpenStudy (gracygirl):

Thank you!